Circular No. 8779

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2006 W4 (HILL) R. E. Hill reports the discovery of a comet in the course of the Catalina

Sky Survey (discovery observation below), the object noted as having a tail ∼ 15′′ long to the west-northwest. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, the object has been found to show cometary appearance by other CCD observers, including G. Hug (Eskridge, KS, 0.7-m reflector, Nov. 23.2 UT; diffuse with a slight coma extension in p.a. ∼ 310^{o}) and P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m reflector, Nov. 24.0; very concentrated, almost-stellar coma of diameter 6′′ with a tail ∼ 6′′ wide and 30′′ long in p.a. 285^{o}).

2006 UT  α_{2000}  δ_{2000}  Mag.
Nov. 22.33647  3^{h}11^{m}58.^{s}56  +26^{o}23′11′′.9  19.3

The available astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2006-W67.

T = 2008 Feb. 29.798 TT  ω = 216.^{o}177
e = 0.29858  Ω = 247.097 }2000.0
q = 4.54345 AU  i = 27.277
a = 6.47755 AU n^{o} = 0.059784 P = 16.5 years

SUPERNOVAE 2006or, 2006os, 2006ot Three type-Ia supernovae have been found on unfiltered CCD images:

2006or in NGC 3891 by T. Puckett and A. Kroes (cf. IAUC 8778), 2006os in UGC 2384 by R. Quimby and F. Castro (cf. IAUC 8754), and 2006ot in ESO 544-G31 by N. Joubert and W. Li (LOSS/KAIT; cf. IAUC 8771).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006or Nov. 18.41 11^{h}48^{m}03.^{s}47 +30^{o}21′24′′.0 16.3 1′′.0 E, 10′′.7 S
2006os Nov. 21.18 2 55 01.01 +16 00 34.8 17.2 6′′.2 E, 10′′.4 S
2006ot Nov. 22.28 2 15 04.84 −20 45 58.2 17.4 3′′.3 E, 5′′.4 N

Additional magnitudes for 2006or: 2005 Dec. 3 UT, [19.1; 2006 Nov. 21.40, 16.3. SN 2006or was ≈ 5 days past maximum on Nov. 22.5 (details on CBET 750). SN 2006os is near maximum brightness (cf. CBET 751). SN 2006ot was not present on a KAIT image taken on Oct. 20.39 (limiting mag 18.5). SN 2006ot is prior to maximum light (details on CBET 754).

2006 November 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green